OpenStudy (anonymous):

solve each proportion. 1. 6/4=x/18 2. 33/b=15/45

OpenStudy (anonymous):

\[\frac{6}{4}=\frac{x}{18}\implies x=\frac{18\times 6}{4}\] then reduce

OpenStudy (anonymous):

1. x=27 2. x=99